### v1

# Intrinsic orbital fourfold anisotropic magnetoresistance in Dirac materials

Abstract
Fourfold anisotropic magnetoresistance (AMR) have been widely observed in quantum materials, but the underlying mechanisms remain poorly understood. Here we find, in a variety of three-dimensional Dirac materials that can be unifiedly described by the massive Dirac equation, the intrinsic orbital magnetic moment of electrons vary synchronously with the magnetic field and give rise to a $\pi$ periodic correction to its velocity, further leading to unusual fourfold AMR, dubbed intrinsic orbital fourfold AMR. Our theory not only explains the observation of fourfold AMR in bismuth but also uncovers the nature of the dominant fourfold AMR in thin films of antiferromagnetic topological insulator $\mathrm{MnBi_{2}Te_{4}}$, which arises from the near cancellation of the twofold AMR from the surface states and bulk states due to distinct spin-momentum lockings. Our work provides a new mechanism for creation and manipulation of intrinsic fourfold AMR in both conventional conductors and various topological insulators.
Introduction.–Anisotropic magnetoresistance (AMR) is a fundamental phenomena in magnetic material, usually arising from the interaction between the electrons and magnetizations, has many useful functionality in magnetic sensors and data recording technologies (Ejsing et al., 2004; Bason et al., 2006; Chappert et al., 2007). Previous studies showed that both AMR and the related planar Hall effect (PHE) widely oscillates in $\pi$ (twofold symmetric component) with respect to the angle between the direction of electric current and magnetic field (Tang et al., 2003). Notably, fourfold ARM was observed in a wide variety of materials (van Gorkom et al., 2001; Muduli et al., 2005; Limmer et al., 2006; Rushforth et al., 2007; Bason et al., 2009; Naftalis et al., 2011; Annadi et al., 2013; Ding et al., 2013; Xiao et al., 2015; Oogane et al., 2018; Ahadi et al., 2019; Zeng et al., 2020; Wadehra et al., 2020; Battilomo et al., 2021; Dai et al., 2022; Song et al., 2022), such as ferromagnetic CoFe alloys (Zeng et al., 2020), FePt epitaxial films (Dai et al., 2022) and antiferromagnetic $\mathrm{EuTi_{2}O_{3}}$ (Ahadi et al., 2019) and $\mathrm{Nd_{2}Ir_{2}O_{7}}$ (Song et al., 2022), which has been attributed to anisotropic relaxation time, higher order perturbation of spin-orbit coupling (SOC) and cluster magnetic multipoles of spins (Yahagi et al., 2020; Dai et al., 2022; Song et al., 2022). In reality, the fourfold component usually appears a subordinate correction and superposes on the profile of twofold part. How to generate and control the predominant fourfold AMR in quantum materials is far unexplored.
Three-dimensional (3D) Dirac materials including the conventional conductors (Zawadzki, 2017; Andreev and Spivak, 2018), the topological insulators (TIs), topological semimetals have attracted much attention due to the fascinating quantum phenomena and the promising applications in low-energy cost electronics, spintronics and plasmonics (Hasan and Kane, 2010; Qi and Zhang, 2011). Dirac materials host Dirac fermions that possess the unique energy dispersion and nontrivial Berry curvature, facilitating the realization of novel topological phase of matter, such as the quantum anomalous Hall effect, axion insulator and the Majorana excitations for promising topological quantum computations (Weng et al., 2015; Chang et al., 2023). It has been shown that the presence of the Berry curvature accompanied with self-rotation of electronic wave-packet, leading to the orbital magnetic moment (OMM) that share the same symmetry properties as the Berry curvature (Xiao et al., 2010). Here, we demonstrate numerous and easily accessible Dirac materials provide a platform to explore novel electromagnetic response associated with intrinsic OMM of electrons. Recently, both bismuth (Yang et al., 2020) and $\mathrm{MnBi_{2}Te_{4}}$ (Wu et al., 2022) that are described by massive Dirac equation exhibit notable fourfold AMR, but the unified mechanism is still lacking.
In this Letter, we show the fourfold AMR originates from the intrinsic OMM of electrons via modifying the velocity in 3D Dirac materials. We could quantitatively explain the fourfold AMR and their anomalous evolution of AMR with the magnetic fields in both bismuth and $\mathrm{MnBi_{2}Te_{4}}$. It has been shown that the twofold AMR of the surface states in TIs could cancel with that of the bulk states due to the distinct spin-momentum locking, leading to a dominant fourfold part. In addition, the fourfold AMR could appear in $\mathrm{Bi_{2-x}Sb_{x}Te_{3}}$ through tuning the carrier density by chemical doping and electrical gating.
Formalism.–To investigate the novel phenomena associated with OMM in 3D Dirac materials within the semiclassical regime, we start from the the semiclassical dynamics of the electronic wave packet, which includes both the Berry curvature and OMM (Sundaram and Niu, 1999; Xiao et al., 2010)
$\dot{\mathbf{r}}=\frac{1}{\hbar}\frac{\partial\bar{\varepsilon}_{n\mathbf{k}}}{\partial\mathbf{k}}-\dot{\mathbf{k}}\times\mathbf{\Omega}_{n\mathbf{k}},\,\hbar\dot{\mathbf{k}}=-e(\mathbf{E}+\dot{\mathbf{r}}\times\mathbf{B}),$ (1)
where $\mathbf{r}$ and $\mathbf{k}$ are the position and momentum of the center of the wave packet of Bloch electrons. It can be seen that, the Berry curvature $\mathbf{\Omega}_{n\mathbf{k}}=\nabla_{\mathbf{k}}\times\left\langle u_{n\mathbf{k}}\right|i\nabla_{\mathbf{k}}\left|u_{n\mathbf{k}}\right\rangle$ gives rise to an anomalous velocity, where $|u^{n}\rangle$ is the periodic part of the Bloch wave function. The OMM of state $\left|u_{n\mathbf{k}}\right\rangle$, $\mathbf{m}_{n\mathbf{k}}=i\left(e/2\hbar\right)\left\langle\nabla_{\mathbf{k}}u_{n\mathbf{k}}\right|\left(\varepsilon_{n\mathbf{k}}-H\left(\mathbf{k}\right)\right)\times\left|\nabla_{\mathbf{k}}u_{n\mathbf{k}}\right\rangle$, modifies the electron energy as $\bar{\varepsilon}_{n\mathbf{k}}=\varepsilon_{n\mathbf{k}}-\mathbf{m}_{n\mathbf{k}}\cdot\mathbf{B}$, where $H\left(\mathbf{k}\right)$ is the Bloch Hamiltonian (Xiao et al., 2005; Thonhauser et al., 2005). The OMM can be understood as an additional magnetic moment caused by self rotating of the wave packet around its center in k space, in contrast to the conventional OMM for the orbital motion driven by Lorentz force.
The electric current density is given by
$\mathbf{J}=-e\intop[d\mathbf{k}]D^{-1}f_{\mathbf{k}}\mathbf{\dot{\mathbf{r}}},$ (2)
where $D=\left[1+\frac{e}{\hbar}(\mathbf{B}\cdot\mathbf{\Omega})\right]^{-1}$ is the modification to the density of states in the phase space. In the presence of the spatially homogenous external fields, the distribution function $f_{\mathbf{k}}$ can be determined by the Boltzmann equation $\dot{\mathbf{k}}\cdot\partial_{\mathbf{k}}f_{\mathbf{k}}=-\frac{f_{\mathbf{k}}-f_{0}}{\tau(\mathbf{k})}$ within the relaxation time approximation (Ziman, 2001), where $\tau(\mathbf{k})$ is the relaxation time and $f_{0,\mathbf{k}}=\left\{\mathrm{exp}\left[\frac{1}{kT}(\bar{\varepsilon}_{\mathbf{k}}-\mu)\right]+1\right\}^{-1}$ is the unperturbed Fermi distribution function. In the linear-response regime, We have $f_{\mathbf{k}}=f_{0,\mathbf{k}}+f_{1,\mathbf{k}}$ with
$f_{1,\mathbf{k}}=-\tau(\mathbf{k})D\left[e\mathbf{E}\cdot\bar{\mathbf{v}}_{\mathbf{k}}+\frac{e^{2}}{\hbar}(\mathbf{E}\cdot\mathbf{B})(\mathbf{\Omega}\cdot\bar{\mathbf{v}}_{\mathbf{k}})\right]\left(-\frac{\partial f_{0}}{\partial\varepsilon}\right).$
with $\bar{\mathbf{v}}_{\mathbf{k}}=\mathbf{v}_{k}-\frac{1}{\hbar}\partial_{\mathbf{k}}(\mathbf{m}_{\mathbf{k}}\cdot\mathbf{B})$ being the velocity of electrons. Substituting $f_{\mathbf{k}}$ and $\dot{\mathbf{r}}$ into Eq. (2) yields the electric conductivity tensor
$\sigma_{\alpha\beta}$ $=e^{2}\intop[d\mathbf{k}]\tau(\mathbf{k})D\left(-\frac{\partial f_{0}}{\partial\varepsilon}\right)\left(\bar{v}_{\alpha}\bar{v}_{\beta}+\frac{e}{\hbar}\bar{v}_{\alpha}B_{\beta}(\mathbf{\Omega}\cdot\bar{\mathbf{v}}_{\mathbf{k}})\right.$
$+\frac{e}{\hbar}\bar{v}_{\beta}B_{\alpha}(\mathbf{\Omega}\cdot\bar{\mathbf{v}}_{\mathbf{k}})\left.+\frac{e^{2}}{\hbar^{2}}B_{\alpha}B_{\beta}(\mathbf{\Omega}\cdot\bar{\mathbf{v}}_{\mathbf{k}})^{2}\right),$ (3)
which allows us to investigate the impacts of OMM on the transport properties of electrons, in particular, beyond the quantitative corrections (Dai et al., 2017; Nandy et al., 2018). Note that we have omitted the terms solely associated with Berry curvature such as the anomalous Hall effect and the chiral magnetic effect of Weyl fermions as well as these higher order terms relevant to the positional shift (Gao et al., 2017; Gao, 2019; Huang et al., 2023).
Massive Dirac fermions.–Dirac equation plays a crucial role in understanding and realizing the fascinating topological phases of matter as well as the novel electromagnetic responses in a large family of quantum materials, such as TIs, topological semimetals) (Shen, 2013; Armitage et al., 2018; Lv et al., 2021). We here utilize the effective Dirac model with a modified mass, which has been used to describe the topological Anderson localization, negative MR and resistivity anomaly in 3D topological materials (Guo et al., 2010; Dai et al., 2017; Fu et al., 2020)
$H_{0}(\mathbf{k})=\left(\begin{array}[]{cccc}\mathcal{M}(\mathbf{k})&0&Ak_{z}&Ak_{-}\\ 0&\mathcal{M}(\mathbf{k})&Ak_{+}&-Ak_{z}\\ Ak_{z}&Ak_{-}&\mathcal{-M}(\mathbf{k})&0\\ Ak_{+}&-Ak_{z}&0&-\mathcal{M}(\mathbf{k})\end{array}\right),$ (4)
where $\mathcal{M}(\mathbf{k})=M-Fk^{2}$ is the gap, and $k_{\pm}=k_{x}\pm ik_{y}$. Both the conduction and valence bands are doubly degenerate $\varepsilon_{\mathbf{k},0,\pm}=\pm\varepsilon$ with $\varepsilon=\sqrt{A^{2}k^{2}+\mathcal{M}^{2}\left(\mathbf{k}\right)}$. The two wave functions of electrons in the conduction bands are $|u_{1,2}(\mathbf{k})\rangle$. Some cumbersome calculations give us the $\mathrm{SU}(2)$ Berry connection $\boldsymbol{\mathscr{\mathcal{A}}}=A^{2}\mathbf{k}\times\boldsymbol{\sigma}/2\varepsilon\left[\varepsilon-\mathcal{M}\left(\mathbf{k}\right)\right]$ and the corresponding Berry curvature
$\varOmega_{i}$ $=\frac{A^{2}}{2\varepsilon^{3}}\left\{(\mathcal{M}+2Fk^{2})\sigma_{i}+\left[A^{2}(\mathcal{M}-\varepsilon-2Fk^{2})\right.\right.$
$\left.\left.+4F\mathcal{M}(\varepsilon-\mathcal{M})\right]\frac{(\mathbf{k}\cdot\boldsymbol{\sigma})k_{i}}{(\varepsilon-\mathcal{M})^{2}}\right\},$ (5)
It can be verified that the OMM and Berry curvature here satisfy the relation of $\mathbf{m}=\frac{e}{\hbar}\varepsilon\mathbf{\varOmega}$ due to the particle-hole symmetry, as the standard Dirac equation (Chang and Niu, 2008).
To reveal the impacts of the magnetic field on transport properties in the planar Hall geometry, we consider a Zeeman term
$H_{z}=\frac{\mu_{B}}{2}\left(\begin{array}[]{cccc}g_{\perp}B_{z}&g_{\parallel}B_{-}&0&0\\ g_{\parallel}B_{+}&-g_{\perp}B_{z}&0&0\\ 0&0&g_{\perp}B_{z}&g_{\parallel}B_{-}\\ 0&0&g_{\parallel}B_{+}&-g_{\perp}B_{z}\end{array}\right),$ (6)
where $B_{\pm}=B_{x}\pm iB_{y}$, $g_{\perp/\parallel}$ are Lande g factor and $\mu_{B}$ is the Bohr magneton. $H_{z}$ breaks the SU(2) symmetry of electrons in the conduction and valance bands as well as the relevant Berry curvature and OMM.
In order to gain clear insights into the impacts of OMM on AMR, we consider an isotropic g-factor $g_{\perp}=g_{\parallel}=g$ and a constant mass $\mathcal{M}\left(\mathbf{k}\right)=M$. The corresponding eigenstates can be solved analytically, and the energy of the top conduction band is modified to $\varepsilon_{\mathbf{k}}(\mathbf{B})=\sqrt{\varepsilon_{k,0}^{2}+\widetilde{B}^{2}+2\widetilde{B}\sqrt{A^{2}(\mathbf{k}\cdot\mathbf{n})^{2}+M^{2}}}$, where $\widetilde{B}=\frac{\mu_{B}}{2}g|\mathbf{B}|$ is the Zeeman energy, $\mathbf{n}=\mathbf{B}/|\mathbf{B}|$ is unit vector along the direction of the magnetic field. For a magnetic field along the $z$ direction, the Berry curvature and OMM of conduction band can be obtained as [details are given in the Supplemental Materials (SM) (SMA, )]
$\mathbf{\Omega}\left(\mathbf{k}\right)$ $=-\frac{\widetilde{B}A^{4}}{2\gamma\varepsilon_{+,\downarrow}^{3}}\left(k_{x}k_{z},k_{y}k_{z},\frac{\gamma(\widetilde{B}^{2}+\gamma)}{\widetilde{B}^{2}A^{2}}\right),$ (7)
with $\gamma=\sqrt{\widetilde{B}^{2}(A^{2}k_{z}^{2}+M^{2})}$. The relevant OMM is given as $\mathbf{m}=\varepsilon_{+,\downarrow}\mathbf{\Omega}$. One finds that the direction of Berry curvature/OMM almost synchronously changes along with the direction of the magnetic field, leading to a nonzero anomalous Hall effect $\sigma_{xy}\neq 0$. More importantly, the correction to the energy dispersion from the OMM $\mathbf{m}_{\mathbf{k}}(\mathbf{B})\cdot\mathbf{B}$ depends on the angle of the magnetic field with a period of $\pi$. As a result, the $\pi/2$-peroidic dependence of electric conductivity on the angle of the magnetic field emerges through the quadratic term like $\left(\mathbf{m}_{\mathbf{k}}\left(\mathbf{B}\right)\cdot\mathbf{B}\right)^{2}$ in Eq. (3), namely, the intrinsic orbital fourfold AMR.
There are several salient features of the intrinsic orbital fourfold AMR here. First, it is even in the magnetic field and odd in the relaxation time, reflecting the magnetoresistance nature (Shtrikman and Thomas, 1965), in contrast to other in-plane magnetotransport phenomena (Cullen et al., 2021; Cao et al., 2023; Wang et al., 2024). Second, this orbital fourfold AMR depends on the form of unique spin-momentum locking of 3D Dirac electrons and the associated intrinsic OMM but does not require the extrinsic anisotropic relaxation time (Dai et al., 2022) or special crystal symmetry for usual ferromagnets (Yahagi et al., 2020). Third, unlike the fourfold AMR in conventional magnetic materials, the permanent magnetic orders from the spin degree of freedom of electrons is not a perquisite for the present fourfold AMR. Next we would like to apply our theory to the observed fourfold AMR in two representative Dirac materials: semimetal bismuth and metallic $\mathrm{MnBi_{2}Te_{4}}$.
AMR in bismuth.–Bismuth, elemental semimetal, exhibits many intriguing quantum phenomena (Gygax et al., 1986; Collaudin et al., 2015; Feldman et al., 2016; Ito et al., 2016; Zhu et al., 2018) and hosts new higher order topological phases (Schindler et al., 2018; Aggarwal et al., 2021). Fig. 1(a) shows that there are three electron pockets near the equivalent $L$-point and hole pocket near the $T$-point (Liu and Allen, 1995). The electronic states near the $L$-point can be effectively described by the Dirac-Wolff Hamiltonian and exhibit small effective mass (about $10^{-3}m_{e}$ with bare electron mass $m_{e}$) and large anisotropic g-factor (several hundreds) (Wolff, 1964; Fuseya et al., 2015), which make it very sensitive to the magnetic field. Recently, the fourfold AMR/PHE has been observed when the electric current and the rotated magnetic field are in the binary-bisectrix plane in micro-thick (111) thin films of single-crystal bismuth (Yang et al., 2020). However, the microscopic mechanism is still under debate (Bi, 4). It was attributed to the anisotropic orbital magnetoresistance together with the chiral anomaly scenario (Yang et al., 2020), while the anomalous angular-dependence in AMR/PHE was ascribed to the field-dependent carrier densities in strong fields (Yamada and Fuseya, 2021).
Figure 1: (a) Schematic illustration of Fermi surface in the binary-bisectrix plane of Bismuth. Comparison between experimental data $\rho_{11}$ of Bismuth (colored scatters) (dat, ) and theoretical results of amplitude ratio of AMR (solid and dashed lines) at 2 K (b) and 300 K (c). The fourfold AMR increases as $B^{2}$ for the weak magnetic field, then decreases as $\left|B\right|^{3}$, and the position of this peak is affected by temperature. The parameters in numerical calculations are given (SMA, ).
Here we numerically calculate the AMR with the OMM, present the results in Figs. 1(b, c) and make some comparison with experimental ones ($\rho_{11}$ from 0.1 T to 2 T in Figs. S5-6 in Ref. (Yang et al., 2020)). As shown in Fig. 1(c), at 300 K, the twofold AMR can be fitted with $a_{1}B$, and the fourfold AMR can be fitted with $a_{2}B^{2}+a_{3}\left|B\right|^{3}$, where $a_{1}$ and $a_{2}$ are positive, but $a_{3}$ is negative. We find that the first two orders of the expansion of $\left[\partial_{\mathbf{k}}(\mathbf{m}_{\mathbf{k}}\left(\mathbf{B}\right)\cdot\mathbf{B})\right]^{2}$ make primary contribution to the fourfold AMR, also support the conclusion via fitting (see Eq. (S17) in SM (SMA, )). It should be noted that the peak in the fourfold AMR at 2 K can be understood from the approximative expension above. Specifically, the positive $a_{2}B^{2}$ term dominates at low field, the negative $a_{3}\left|B\right|^{3}$ term becomes important as increase the magnetic field, inevitably producing the maximum. The fourfold AMR at 2 K and 300 K primely lie in the decreasing regime and the increasing regimes due to the thermal broadening, respectively. The calculations are in quantitative agreement with the experimental results of the existence of the fourfold AMR and the nontrival field and temperature dependences.
Roles of surface states.–As the band inversion occurs, the material would enter the topological insulating phase, supporting two-dimensional (2D) Dirac surface states (Bansil et al., 2016). The effective model for the surface states in the $x-y$ plane can be written as (Zhang et al., 2009; Shan et al., 2010; Jiang et al., 2022)
$H_{sur}=\hbar v_{F}(k_{x}\sigma_{y}-k_{y}\sigma_{x}),$ (8)
which resembes the 2D Rashba SOC $\left(\boldsymbol{k}\times\boldsymbol{\sigma}\right)\cdot z$ (Manchon et al., 2015). Here $v_{F}$ is the effective Fermi velocity of the surface states, $\sigma_{x,y}$ are the Pauli matrices for the real spin of electrons. Note the Zeeman term of an in-plane magnetic field, $H_{z}=\frac{1}{2}g\mu_{B}(B_{x}\sigma_{x}+B_{y}\sigma_{y})$, does not open a gap but only shifts the position of surface Dirac point (Zhang et al., 2021). It should be pointed out that, for 2D systems, the OMM is usually normal to the plane and would not modify the energy dispersion and velocity of electrons.
Figure 2: (a) The twofold AMR of the 2D surface states and the bulk states in antiferromagnetic TI $\mathrm{MnBi_{2}Te_{4}}$ (Wu et al., 2022), which amplitudes have opposite sign. (b) Evolution of AMR with the magnetic field. As the magnetic field increases, the fourfold AMR becomes noticeable and then dominates over the twofold part around 10 T. In our calculations, the Fermi energy $E_{F}=500\;\mathrm{meV}$ is measured from the center of the bulk band gap and T= 10 K. Other parameters are obtained from first-principle calculations (MBT, ).
To simulate the scatterings between surface Dirac electrons and the localized magnetic atoms (such as Mn atoms in $\mathrm{MnBi_{2}Te_{4}}$ (Wu et al., 2022)), we consider the spin-dependent scattering potential as $U(r)=\mu_{0}\sum_{i}\mathbf{M}\cdot\mathbf{\boldsymbol{\mu}}\delta(\mathbf{r}-\mathbf{R}_{i})$ (Chiba et al., 2017), where $\mu_{0}$ is the interaction strength, $\mathbf{\boldsymbol{\mu}}=\frac{1}{2}g\mu_{B}\mathbf{\boldsymbol{\sigma}}$ is spin magnetic moment of electrons, $R_{i}$ is the impurity position, and $\mathbf{M}$ is the field-polarized magnetic moment of the local impurities. We can calculate the relaxation time in the Born approximation and the electric conductivity of the surface states (SMA, ). The calculated conductivities for the surface and bulk states are shown in Fig. 2(a). The twofold AMR of the surface states can be understood from the distinction of the spin-momentum lockings. For the Dirac surface electrons, when the spins of electrons are parallel or antiparallel to the magnetic field or induced magnetization, the backscattering is forbidden, the corresponding resistivity will be much less affected by the magnetic field ($\rho_{\perp}$). On the other hand, when the spins are perpendicular to the field, the backscattering becomes allowed due to the broken time reversal symmetry, resulting in a significant enhancement of resistivity ($\rho_{\sslash}$). That is a positive anisotropic resistivity (i.e., $\rho_{\sslash}-\rho_{\perp}>0$). However, for the bulk states with parallel spin and momentum $\boldsymbol{k}\cdot\boldsymbol{\sigma}$, the direction of field enhanced resistivity and the direction of resistivity that is less sensitivitive to magnetic fields get exchanged, leading to a negative anisotropic resistivity ($\rho_{\sslash}-\rho_{\perp}<0$). It was also supported by the specific calculations of the anisotropic resistivity of a 2D slice of bulk states (SMA, )
$\sigma_{xx}^{2D}=\frac{8e^{2}\hbar v_{F}^{2}\left[1+(2+\sqrt{3})\cos 2\theta\right]}{\sqrt{3}n_{s}(\mu_{0}g\mu_{B})^{2}|\mathbf{M}|^{2}},$ (9)
where $n_{s}$ is the density of magnetic impurities and $\theta=\mathrm{arctan}\;(M_{y}/M_{x})$. One can find that the cancellation of the twofold AMR between the surface states and the bulk states probably stems from the distinct spin-momentum lockings. In fact, this cancellation mechanism should be generally applicable to a wide range of 3D TIs including nonmagnetic, magnetic and higher-order ones (Xu et al., 2019; Zhang et al., 2019a, 2020).
When the Fermi level lies in the conduction band of bulk, both the surface states and the bulk states simultaneously contribute to the electric transport. Because of the much hihger mobility, the surface states dominates over the bulk ones at lower fields. It is probable that, under some proper conditions, they can almost cancel out, making the fourfold AMR dominant. We numerically calculate the AMR that from both the bulk and the surface states of $\mathrm{MnBi_{2}Te_{4}}$ which is shown in Fig. 2(b). It can be seen that when the magnetic field is around 10 T, the fourfold AMR dominates. Further increasing of the magnetic field, the twofold AMR becomes dominant again, but the sign of its amplitude changes, which is in line with the observed AMR (Wu et al., 2022).
Recently, the twofold AMR of Dirac surface states has been observed in the dual-gated devices of thin $\mathrm{Bi}{}_{2}\mathrm{SbTe}_{3}$ (Taskin et al., 2016; Zheng et al., 2020), where the Fermi level lies in the bulk gap. The electrical gating and chemical doping could tune the carrier density and shift the Fermi level into the bulk energy band, facilitating the realization of the predominant fourfold AMR therein.
Summary.–We have found the tunable intrinsic orbital fourfold AMR due to OMM of electrons in various 3D Dirac materials. The unique spin-momentum lockings in topological bulk states and surface states cause competition between their twofold AMR, leaving predominant fourfold component. Our calculations are in quantitative agreement with the fourfold AMR in both bismuth and $\mathrm{MnBi_{2}Te_{4}}$ and the nontrivial magnetic-field dependence. This work reveals the significance of intrinsic OMM in novel AMR, inspiring more investigations of intriguing quantum phenomena associated with OMM in Dirac materials, such as the nonlinear PHE (He et al., 2019). Moreover, the first-principles calculations of OMM in the semiclassical transport theory could provide us a new and significant ingredient to deep investigate the novel magnetotransport properties of realistic quantum materials with complicated Fermi surfaces (Zhang et al., 2019b) and relaxation time (Dai et al., 2022).
The authors acknowledge Y. Gao, Q. Niu, T. Qin and Z. Yuan for insightful discussions. This work was supported by the National Natural Science Foundation of China under Grants (No. U2032164 and No. 12174394) and by the High Magnetic Field Laboratory of Anhui Province and by the HFIPS Director’s Fund (Grant Nos. YZJJQY202304 and BJPY2023B05) and Chinese Academy of Sciences under contract No. JZHKYPT-2021-08.